A factory manufactures cars with a warranty of 5 years on the engine and transmission.  An engineer believes that the engine or transmission will malfunction in less than 5 years.  He tests a sample of 40 cars and find the average time to be 4.8 years with a standard deviation of 0.50. At a 2% confidence level, is there enough evidence to support the idea that the warranty should be revised ?  ( P-value Test)


1
Expert's answer
2021-02-24T07:31:36-0500

The provided sample mean is "\bar{x}=4.8" and the sample standard deviation is "s=0.50."

The sample size is "n=40." The number of degrees of freedom are "df=n-1=39," and the significance level is "\alpha=0.02."

The following null and alternative hypotheses need to be tested:

"H_0: \mu\geq5"

"H_1: \mu<5"

This corresponds to left-tailed test, for which a t-test for one mean, with unknown population standard deviation will be used.

Based on the information provided, the critical value for a left-tailed test is "t_c=-2.124742."

The rejection region for this two-tailed test is "R=\{t:t<-2.124742\}"

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{4.8-5}{0.5/\sqrt{40}}=-2.529822"

Since it is observed that "t=-2.529822<-2.124742=t_c," it is then concluded that the null hypothesis is rejected. Therefore, there is enough evidence to claim that the population mean "\mu" is less than 5, at the 0.02 significance level.

Using the P-value approach: The p-value "df=39, t=-2.529822," "left-tailed," is "p=0.00778210," and since "p=0.00778210<0.02," it is concluded that the null hypothesis is rejected. Therefore, there is enough evidence to claim that the population mean "\mu" is less than 5, at the 0.02 significance level.


Therefore there is enough evidence to support the idea that the warranty should be revised at the 0.02 significance level.